OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 34 inches. Find the length and width of the rectangle

OpenStudy (anonymous):

Let w be the rectangle's width. Solve the following for w: 2 (w + (w + 4)) = 34 w = 13/2

OpenStudy (anonymous):

\[\text{Length}=\frac{13}{2}+4=\frac{21}{2} \]

OpenStudy (anonymous):

 Heres something a little bit simpler and maybe easier to follow. Finding the perimeter is finding the sum of all the sides. The width=w, and the length=w+4. There are two of each of these, therefore, you have to add them all. w+w=2w, (w+4)+(w+4)=2w+8. 2w+2w+8= 4w+8. Set this expression equal to 34, the perimeter of the rectangle, and solve for w. By doing that, you should get w=6.5. Which is equal to (13/5). To find the length of the rectangle, simply plug 6.5 into w+4. 6.5+4=10.5 The length of the rectangle is 10.5 inches.